A Ferris wheel is moving at an initial angular velocity of 1.0 rev/43 s. If the operator then brings it to a stop in 2.6 min, what is the angular acceleration of the Ferris wheel? Express your answer in rad/s2. rad/s2

Through how many revolutions will the Ferris wheel move while coming to a stop?

please explain!!!
Answer #1

w1 = 2*pi/43 = 0.146 rad/s

w2 = 0


a) alfa = (w2-w1)/t

= (0-0.146)/(2.6*60)

= -9.362*10^-4 rad/s^2


b) theta = w1*t + 0.5*alfa*t^2

= 0.146*2.6*60 - 0.5*9.362*10^-4*(2.6*60)^2

= 11.38 rad

= 1.81 revolution